Assume that your father is now 55 years old and plans to retire after 5 years from now.

He is expected to live for another 15 years after retirement. He wants a fixed retirement

income of Rs. 1,00,000 per annum. His retirement income will begin the day he retires,

5 years from today, and then he will get 14 additional payments annually. He expects to

earn a return on his savings @ 10% p.a., annually compounding. How much (to the

nearest of rupee) must your father save today to meet his retirement goal?


1
Expert's answer
2021-05-26T16:26:01-0400

P = Amount required annually = 100000

n = 15 years

r = return = 10%

Amount required at retirement "= P + \frac{P \times (1 - (1+r)^{-(n-1)}) }{ r}"

"= 100000 + \frac{100000 \times (1 - (1+0.1)^{-(15-1)}} {0.1} \\

= 100000 + \frac{100000 \times 0.736687545}{ 0.1} \\

= 100000 + 736668.7457

=836668.7457"

The amount required at retirement is =836668.7457

Calculation of Annual savings:

n = 5 years

r = annual return = 10%

Let P = Annual Savings required

P*((1+r)n-1)/r

​836668.7457=p*(1+0.1)5-1/0.1

"0.61051p=83668.7457"

"P=137047.2977"